### v3

# organization=Instituto Federal de Educação Ciência e Tecnologia do Ceará (IFCE), Campus Tianguá,addressline=Av Tabelião Luiz Nogueira, s/n - Santo Antônio, city=Tianguá, postcode=62324-075, state=Ceará, country=Brazil

Abstract
In this letter, we propose a relationship between the so-called Hubble-Lemaître constant $H_{0}$ and holographic complexity related to the emergence of spacetime in quantum gravity. Such a result can represent an important step to understanding the Hubble tension by introducing a quantum gravity perspective for cosmological observations: regarding the degree of quantum complexity we measure around us.
journal: Physics Letters B
1 Introduction
In recent years, a contention has arisen in cosmology, the so-called Hubble tension [1, 2, 3]. Such a problem has its origin in a controversy between early-universe measurements of the so-called Hubble-Lemaître constant $H_{0}$ provided by the Planck collaboration [4], and late-time observations provided by SHOES [5]. Such a discrepancy has been confirmed by several other late-time observations [5, 6, 7, 8, 9, 10, 11, 12], which have shown a clear distinction from the value provided by the Planck satellite.
The early-universe result is based on the analysis of the Cosmic Microwave Background radiation (CMB) which was formed when the universe was in its childhood. On the other hand, the late-time universe measurements are based on several observations related to objects formed in epochs when the universe was more old and complex such as cepheids, quasars, red giants stars, etc.
Putting it in numbers, the early universe result has provided a value of $H_{0}=67.4\pm 0.5\;km\;s^{-1}Mpc^{-1}$. On the other hand, the late-time universe measurements have provided values for $H_{0}\geq 70$. For example, the latest SHOES measurement gives $H_{0}=73.2\pm 1.3\;km\;s^{-1}Mpc^{-1}$ [12].
In statistical terms, the difference between early and late-time results amounts to $4\sigma$ to $6\sigma$ depending on the datasets considered [2]. Such a discrepancy has cast doubt on the very constancy of $H_{0}$, and is considered sufficient to establish a crisis in cosmology, asking for new physics [13, 14, 15, 16, 17, 18, 19].
Several proposals have been suggested to mend the $\Lambda$CDM model and explain the different values measured to $H_{0}$. An extensive list of such proposals can be found in [2]. In front of these efforts, it has been argued that a deeper breakdown of our understanding of the cosmos becomes necessary to solve the Hubble tension.
A forthcoming quantum version of cosmology, based on a still unknown theory of quantum gravity, can promote such a paradigm shift, by reformulating our view about the nature of spacetime itself. However, so far, no satisfactory proposal to solve the Hubble tension has been introduced by any candidate theory to explain the quantum aspect of the gravitational phenomena.
In the present letter, we propose a possible explanation for the origin of the Hubble tension by considering new results on quantum gravity introduced recently in [20, 21]. In such a context, we shall find out that $H_{0}$ can be related to the quantum complexity associated with the emergence of spacetime.
Consequently, the discrepancies among the $H_{0}$ values, when one considers the early and late-time measurements, can rely on a natural aspect of the cosmos: the evolution of its complexity throughout time according to the second law of holographic complexity [22]. Such a result may not only shed light on the Hubble tension but also introduce a new perspective for cosmological observations: in terms of holographic quantum complexity we measure around us.
The paper is organized as follows: in sections (2) and (3), we review the main aspects of the theory introduced in [20, 21]. In section (4), we address some basic concepts related to holographic quantum complexity. In section (5), we demonstrate that $H_{0}$ can be related to holographic complexity. Section (6) is devoted to conclusions and discussions.
2 The holographic principle as a prelude to the emergence of spacetime
A forthcoming quantum theory of gravity must provide a radical change in our comprehension of the world. However, possibly the most disruptive lesson quantum gravity must bring us is that related to the role of spacetime in our description of physical reality. In this sense, the holographic principle has been pointed as a possible guide [23, 24, 25].
The most successful form of the holographic principle consists of the AdS/CFT correspondence which consists of a non-perturbative approach to superstring theory [26]. In its original formulation, the AdS/CFT correspondence establishes a duality between a type IIB string theory living in a $AdS_{5}\times S^{5}$ bulk and a conformal field theory (CFT) living on the bulk boundary.
From the necessity of aligning the AdS/CFT ideas with the gravitational dynamics present in our four-dimensional universe, it has been considered its complementarity with the Randall-Sundrum II (RSII) scenario [27, 28, 29, 30]. By taking this way, our four-dimensional universe, which is viewed in the AdS/CFT framework as localized on the AdS boundary, turns to be described by a stack of coincident Randall-Sundrum II branes, and the CFT must be considered coupled to gravity.
Such a complementarity between the AdS/CFT and the Randall-Sundrum frameworks has given birth to a new version of the AdS/CFT correspondence, which has been proposed by Emparan, Kaloper, and Fabri [31], and independently by Tanaka [32] in 2002. It was named by Gregory as the Black Hole Holographic Conjecture (BHHC) [33, 34].
This new version of the AdS/CFT correspondence proposes that the gravitational theory living in the AdS bulk, described by a type IIB string theory, must be holographically related not only to a CFT but to a semiclassical gravitational theory (gravity + quantum corrections) living on the AdS boundary. In such a context, bulk higher-dimensional effects must be translated as quantum corrections to the boundary gravitational theory.
It has opened up the possibility of interesting investigations related to the nature of the boundary theory in the AdS/CFT framework. For example, by applying the BHHC ideas to a cosmological scenario, it has been demonstrated that quantum-corrected Friedmann equations similar to Loop Quantum Cosmology (LQC) ones can be induced on the AdS boundary [20].
Due to the fact that quantum corrections arising in such a context must correspond to the holographic dual of higher-dimensional effects in the AdS bulk, such a result has raised the interesting possibility of connecting the superstring theory in the bulk to a quantum gravitational theory living on the AdS boundary.
In fact, it has been demonstrated that LQG polymer quantization can be applied in such a context. In this case, the quantum degrees of freedom belonging to the boundary gravitational theory correspond to $U(1)$ quantum holonomies which give rise to quantum geometrical states $\psi(x)=e^{ipx}$ defined on the AdS boundary in the following way:
$\hat{h}_{\pm}\psi(x)=e^{ip\sqrt{\Delta}}\psi(x)=e^{ip(x+\sqrt{\Delta})}=\psi(x+\sqrt{\Delta}).$ (1)
Even though such holonomies are similar to that we have in LQG to describe a quantum spacetime, in the scenario introduced in [20], the quantization parameter defining such objects is related to the brane tension and, consequently, to the closed string coupling [35, 36].
In fact, the quantity $\Delta$ in Eq. (1) correspond to a quantum of area parameter given by [20]:
$\Delta=\frac{12\pi}{\sigma}=96\pi^{4}\alpha^{\prime 2}g_{s}=\frac{24\pi^{3}L^{4}}{N}\;,$ (2)
where $\sigma$ is the brane tension, $\alpha^{\prime}$ is the Regge slope, $g_{s}$ is the closed string coupling, $L$ is the AdS curvature radius, and $N$ corresponds to the number of branes sourcing gravitationally the bulk geometry [35, 36].
Such results introduce a new scenario where braneworld geometrical degrees of freedom can be associated with a polymer structure, defined by a graph in the form of a regular lattice:
$\gamma_{\sqrt{\Delta}}=\{x\in\mathbb{R}\mid x=n\sqrt{\Delta},\forall n\in\mathbb{Z}\}\;.$ (3)
It establishes a reinterpretation of branes as $U(1)$ quantum polymer structures. In this way, a possible conection between the main approaches to quantum gravity nowadays is traced, by connecting fundamental objects in string theory and LQG: branes and polymer structures.
Such a result has been useful in addressing interesting problems as those related to the big bang singularity in $AdS/CFT$ correspondence [37, 38], and the Immirzi ambiguity in LQG [39]. However, possible the main lesson it can bring us is that the way branes will be combined to source gravitationally the AdS geometry will depend on the rules governing quantum polymer structures similar to that appear in LQG.
However, at this point, some care needs to be taken since a decision must be made regarding the interpretation of using LQG techniques in the context of the AdS/CFT correspondence. In this sense, the existing LQG literature offers two possibilities.
The first one is that the use of LQG techniques will provide the quantization of the AdS boundary geometry. However, it has been demonstrated that the LQG quantization techniques not only quantize spacetime but also modify it [40]. In this case, one should not have an AdS spacetime anymore after the application of the LQG quantization methods, and the bridge between LQG and string theory would be broken.
The second, and most radical possibility, is that the LQG quantization procedure would not only quantize spacetime, but it will dissolve it into a pre-geometric structure that lives beyond spacetime itself, and from which spacetime ought to emerge.
This is in tune with some AdS/CFT ideas that propose that spacetime must be an emergent aspect of physical reality. In fact, the AdS/CFT correspondence provides the most well-understood example for the emergence of spacetime by proposing that the AdS bulk must be emergent from the dynamics of quantum degrees of freedom living on the AdS boundary [41, 42, 43, 44, 45, 46]. However, the fundamental aspects of how the bulk spacetime, and its gravitational dynamics emerge from the dual theory, are still shrouded in many mysteries.
In this way, the theory introduced in [20] can bring us new insights on what seems to be in the heart of the $AdS/CFT$ correspondence: that spacetime must not be fundamental but ought to emerge from quantum information.
3 A new holographic prescription to the emergence of spacetime.
The emergent nature of spacetime has been proved to be a peculiar feature of quantum gravity. In such a context, it has been proposed that spacetime must emerge from quantum information codified in a set of quantum reference frames connected by quantum correlations [47, 48, 49].
The notion of quantum reference frame arises in the context of quantum mechanics as related to the universality of the theory: both observers and observed must satisfy the same natural laws. In this case, if the quantum description of the world is universal, one would expect that reference frames (observers) should obey the laws of quantum mechanics, corresponding in this way to quantum systems, such as quantum particles.
Considering the string theoretical approach to quantum gravity, a brane is a physical object that generalizes the notion of a point particle. In this way, one can naturally suggest that branes must assume the role of quantum reference frames in such a context.
Such a concern agrees with the result found out in [20] that branes correspond to $U(1)$ holonomies structures. In fact, this result allows us to conceive branes as quantum clocks, i.e., temporal quantum reference frames in a quantum gravity scenario.
However, at this point, a peculiar detail related to the results found out in [20] arises: the holonomy structure of branes can be used not only to establish these objects as quantum reference frames but also to weave the entanglement network between them.
Such a conclusion is rooted in the fact that quantum holonomies $\hat{h}_{ij}$ consist of unitary transformations connecting two quantum geometrical states, belonging (in general) to two different Hilbert spaces, through entanglement [50, 51, 52]:
$\hat{h}_{ij}\;:\;\mathcal{H}_{i}\rightarrow\mathcal{H}_{j}\;,$ (4)
where
$\hat{h}_{ij}\hat{h}_{ij}^{\dagger}=\mathbb{I}\;.$ (5)
In this way, if we label the $\mathcal{H}_{i}$s as the Hilbert spaces of $N$ branes, in the equations above, the indices $i,j$ will range as $i,j=1,...,N$, and the holonomies $\hat{h}_{ij}$ will be promoted to $U(N)$ unitary matrices.
However, as highlighted in [21], the important detail is that there is no fundamental distinction between the holonomies describing the branes serving as quantum reference frames and the holonomies establishing the quantum correlations between them. Note that, in such a construction, one could arbitrarily label some holonomies to correspond to that frozen to define the brane quantum reference frames, while taking the others to correspond to the connections among them. Consequently, even what has been called a quantum reference frame can be understood fundamentally as a quantum correlation, and vice-versa.
In this way, it is possible to introduce a structural realism [53] interpretation to quantum gravity in the context of the results found out in [20, 21], where only quantum correlations are considered as fundamental, and spacetime would be seen as completely emergent from the combinatorial rules governing the structure of such correlations.
Such a perspective can be useful to understand some of the most fundamental aspects related to the emergence of spacetime in quantum gravity. For example, according to the usual form of the holographic principle, spacetime geometry must emerge holographically from a quantum theory living in one spatial dimension lower. But a conceptual problem has been pointed out in such a construction: in the situation where the bulk emerges from the boundary, the last should be more fundamental than the first. It is difficult to reconcile with the fact that, in the AdS/CFT scenario, the duality between the boundary and the bulk theories is exact and thus symmetrical [54, 55, 56, 57].
This kind of conceptual barrier is rooted in the fact that it is a challenge to completely leave the idea of the fundamental existence of spacetime in physics. In fact, how could physics exist beyond spacetime, and how could things exist, and become entangled, for example, without some loci where and when they happen and change? In this case, in the context of the holographic principle, we still need the existence of a locus (the boundary) from which the bulk will emerge.
On the other hand, if only quantum correlations matter, as suggested in [21], combinatorial information can exist without the necessity of a locus for the existence of objects. In fact, in such a scenario, localized objects do not exist anyway, but only correlations without objects correlated must correspond to the elements of physical reality.
It is possible to take a deeper look at the nature of the structure of quantum correlations from which spacetime must emerge by considering an interesting fact pointed out by Girelli and Livine in [47]. According to this fact, a vector in the representation space of the unitary group corresponds to an intertwiner with $N$ legs dressed by $SU(2)$ representations.
In this way, the $U(N)$ group elements $\hat{h}_{ij}$ in the Eqs. (4) and (5) can be written in terms of basis vectors in the representation space as
$\hat{h}_{ij}=e^{M_{ij}}=\mathbb{I}+M_{ij}+\frac{1}{2}M_{ij}M_{ik}+...\;\;\;,$ (6)
where $M_{ij}$ corresponds to a $N$-leg intertwiner belonging to the representation space of $U(N)$. Such an intertwiner corresponds to an abstract spin network.
Consequently, the $U(N)$ quantum states introduced in [21] can be described by a superposition of intertwiner states, and the geometry of the emergent AdS spacetime can be determined by the combinatorial rules governing abstract spin networks.
The results found out in [20, 21] make it possible to set out a new prescription for the emergence of spacetime in the context of the $AdS/CFT$ correspondence as follows:
A type IIB string theory living in a classical $AdS$ bulk emerges from abstract spin networks describing quantum correlations without correlata, in the limit where the number of correlations becomes large.
Important detail here is that abstract spin networks are graphs labeled by $SU(2)$ representations, in the same way we have in full LQG. However, such objects carry only combinatorial information, without any reference to a background geometry or topology. In this way, it is possible to say that, according to the scenario introduced in [21], spacetime must be seen as completely emergent from algebra.
We must highlight that the pre-geometrical regime from which spacetime ought to emerge must occur far from the limit where the number of branes sourcing the bulk geometry is large since, in such a limit, classical spacetime must emerge, according to the $AdS/CFT$ correspondence. It must correspond also to the limit where the number of quantum correlations becomes large.
Moreover, from the coincidence between the number of branes and the number of intertwiners external legs in the approach introduced in [21], the semiclassical limit proposed by the $AdS/CFT$ correspondence aligns with the reasoning introduced in [47] about the semiclassical limit of a pre-geometrical $U(N)$ matrix theory written in terms of $N$-leg intertwiners. In this context, the semiclassical limit is obtained when the number of intertwiners legs $N$ becomes large.
We note that in the context of the theory introduced in [20, 21], the concept of quantum complexity would be useful and could be seen as an intrinsic property of the structure of quantum correlations from which spacetime ought to emerge, being related to patterns of connectivity and transformation within such a structure.
In the next section, we shall explore the concept of holographic complexity in the context of the holographic principle. Moreover, we shall introduce the law governing its evolution.
4 A new window for the emergence of spacetime in quantum gravity: the second law of quantum complexity.
In the context of quantum information theory, the concept of quantum complexity has gained increasing importance in recent times, presenting capillarity in several branches of knowledge [58, 59, 60, 61].
Briefly, complexity expresses the hardness of some task in terms of the amount of computational resources needed to accomplish it. In quantum information theory, any quantum state has an associated complexity, and preparing it using unitary operations is a task in itself. In this context, quantum entanglement has been pointed as the key resource to computational operations [62, 63].
Among the several applications of the concept of quantum complexity, it has been pointed out that it must have a fundamental role in holography. As the holographic principle has been considered a guide to quantum gravity, quantum complexity might be a useful way to understand important aspects of the quantum nature of gravitational phenomena.
Considering the AdS/CFT formulation of the holographic principle, it has been suggested that the complexity associated with the quantum entanglement structure of conformal quantum fields on the boundary of an AdS space might give us insight into the emergence of the geometrical properties of the bulk. [64].
The origin of such an argument is the fact that entanglement alone is not enough to explain how complex bulk geometries emerge [65, 66]. However, by connecting the bulk spacetime properties to the complexity of quantum states, it is possible to trace an alternative way to understand how larger and more intricate bulk geometries can emerge.
Such an idea can be captured by the two main proposals to calculate quantum complexity in the context of the AdS/CFT correspondence, the so-called "Complexity equals to action" [67], and the "Complexity equals to volume conjectures [65, 66].
The first proposal says that the complexity of a quantum state on the AdS boundary must be related to the gravitational dynamics in the bulk spacetime, being proportional to the action in the Wheeler-DeWitt patch. The second proposal suggests that quantum complexity must be related to a volume in the AdS bulk, being proportional to the volume of the maximal codimension-1 bulk slice.
Concerning the second proposal, which will be implemented in the present letter, the holographic complexity $C_{h}$ related to some region on the AdS boundary is given by [64]:
$C_{h}=\frac{Vol(\gamma)}{8\pi LG}=\frac{L^{d-1}}{8(d-1)\pi}\frac{Vol(R)}{\epsilon^{d-1}},$ (7)
where $L$ is the AdS radius, $\gamma$ is the Ryu-Takayanagi (RT) surface [41, 42], and $Vol(\gamma)$ the volume under such a surface in the AdS bulk. The parameter $\epsilon$ denotes a necessary UV cutoff length scale in the boundary theory, and $Vol(R)$ denotes the volume of the boundary spacelike subregion, with a radial size $R$, which goes as $R^{d-1}$. Here $d$ gives the number of bulk spatial dimensions.
Considering the complexity measure above, if quantum complexity increases, bulk geometry also expands. It reinforces the idea that spacetime must be fundamentally linked to quantum computation, being not fundamental but emergent from quantum information.
An important aspect in such discussions is that quantum complexity must obey a second law analogous to the thermodynamical one for entropy [22, 68, 69]. According to such a law: If the computational complexity is less than the maximum, then with the overwhelming likelihood it will increase, till it saturates at a maximal value.
The second law of quantum complexity must have a crucial role in the issue of the emergence of spacetime since it must dictate the way spacetime must emerge from quantum information. In fact, even more implying that spacetime consists of an emergent entity, the second law of quantum complexity must imply that it must emerge in a deterministic way based on the rules of quantum computation.
It is also interesting to point out that the second law of quantum complexity must have a crucial role in the investigation of the evolution of the cosmos due to its relationship with the second law of thermodynamics. For example, it has been argued that the cosmic accelerated expansion can be understood in terms of the universe’s quantum complexity growth with time [70].
In the next section, we shall propose that the evolution of the universe’s holographic complexity by its second law can shed some light on the issue of Hubble tension if one considers the perspective of spacetime as an emergent entity according to the lines introduced in [20, 21].
5 Holographic complexity and the Hubble tension: a possible quantum gravitational standpoint for cosmological observations.
Considering the fundamental aspect of the theory introduced in [20, 21], that the degrees of freedom to the emergence of spacetime consist of quantum correlations only, the concept of quantum complexity must be helpful.
In such a situation, quantum states and unitary transformations would not be thought of as "things" but as patterns of connectivity and transformation within such a structure. Moreover, the growth of complexity would reflect structural change, rather than mere dynamical evolution. In such a context, one would have a way to think about change in a scenario where time is absent.
Taking into account the issue of the emergence of spacetime, one can consider that quantum entanglement carried by the structure of correlations can be used as the key resource for the quantum computational process driving such a phenomenon [62, 63].
In this situation, quantum complexity can be given in terms of the number of quantum correlations effectively used to accomplish spacetime emergence. To obtain such a number, one must fix at first the total number of degrees of freedom in the theory by considering both its UV and IR regimes [71].
Concerning the IR regime, one must put the system into a region with a finite size. In the present scenario, such a region must be large enough to contain the necessary holographic information for the emergence of spacetime.
Particularly, in the $AdS/CFT$ framework, causal diamonds defined on the $AdS$ boundary carry such a necessary amount of information [72, 73, 74], being considered as the natural unit of holography in a region of spacetime [75].
Based on such a fact, we shall take the universe’s causal diamond as the region of finite size used to regularize the IR regime of the present theory for the emergence of spacetime.
In this case, we have that the radius of the universe causal diamond is given by
$R_{\Diamond}$ $=$ $\frac{1}{H_{0}}\int^{\infty}_{-1}\frac{dy}{\sqrt{\Omega_{m0}(1+y)^{3}+\Omega_{\Lambda 0}}}\sim\frac{1}{H_{0}}\;,$ (8)
where $\Omega_{m0}$ and $\Omega_{\Lambda 0}$ correspond to the matter and cosmological constant density parameters, respectively, whose values have been fixed by cosmological observations [4].
On the other hand, to regularize the UV regime, one must introduce a cutoff length scale in the theory. In the scenario introduced in [20, 21], it can be provided by the area parameter $\Delta$ defined by the Eq. (2). In this case, one can take $\epsilon=\Delta^{1/2}$ as the necessary UV cutoff.
Bringing all these ingredients together, the holographic complexity necessary to accomplish the emergence of cosmic spacetime can be written as
$C_{h}\sim L^{3}\frac{R_{\Diamond}^{3}}{(\Delta^{1/2})^{3}},$ (9)
which corresponds to the holographic complexity (7) associated with the universe’s causal diamond, for the case of $d=4$, and a UV cutoff length scale $\epsilon=\Delta^{1/2}$.
Considering the relationship between $C_{h}$ and the number of quantum correlations effectively used to the emergence of cosmological spacetime, we shall take:
$C_{h}\sim N^{\alpha}.$ (10)
In the expression above, the exponent $\alpha$ will provide a measure of how hard the task of picking quantum correlations up to accomplish the emergence of spacetime (as $\alpha$ increases, harder becomes the task). As we shall see, the exponent $\alpha$ will also provide the saturation limit of $C_{h}$.
Taking into account the spin network description of the theory introduced in [21], we shall consider that the number $N$ in the Eq. (10) can be written in terms of the amount of spin network external legs puncturing the universe causal diamond horizon [21].
In this case, we shall consider that the spin network’s legs patch quanta of area on the causal diamond horizon in a fractal way, according to the following relation:
$N\sim\Big{(}\frac{A_{\Diamond}}{\Delta}\Big{)}^{\varepsilon},$ (11)
where $A_{\Diamond}=4\pi R_{\Diamond}^{2}$ is the horizon’s area. In the Eq. above, $\varepsilon>0$ can be related to the fractal dimension attributed to the causal diamond horizon by considering its microscopic structure description [76, 77] (See Fig. (1)).
Figure 1: Artistic representation of spin network legs crossing the universe’s causal diamond horizon. Each spin network leg patches a quantum of area which is distributed on the horizon surface in a fractal way, according to Eq. (11). The number of spin network legs crossing the horizon $N$ can also be related to the amount of quantum complexity necessary for the emergence of cosmic spacetime. (Figure generated by CANVA®  with the use of AI.)
By tracing the basic assumptions of our construction, let us now present our results. In this way, from Eqs. (8) and (11), we can find that
$R_{\Diamond}^{2}H_{0}^{2}$ $\sim$ constant
$N$ $\sim$ $\Big{(}\frac{R_{\Diamond}^{2}}{\Delta}\Big{)}^{\varepsilon}.$ (12)
Such results together provide
$N^{1/\varepsilon}\Delta H_{0}^{2}\sim\textrm{constant},$ (13)
which together with Eq. (2) gives
$N^{\frac{(1-\varepsilon)}{\varepsilon}}L^{4}H_{0}^{2}\sim\textrm{constant}.$ (14)
Now, by taking into account Eqs. (9) and (11), we obtain
$L\sim\frac{C_{h}^{1/3}}{N^{\frac{1}{2\varepsilon}}}.$ (15)
By inserting such result in Eq. (14), one obtains
$H_{0}^{2}\sim N^{\frac{1+\varepsilon}{\varepsilon}}C_{h}^{-4/3}.$ (16)
Finally, by using Eqs. (9) and (16), we obtain
$H_{0}^{2}\sim C_{h}^{\gamma}\;.$ (17)
In this way, $H_{0}$ depends on the holographic complexity related to the emergence of cosmological spacetime through an exponent $\gamma$ given by
$\gamma=\frac{3+3\varepsilon-4\alpha\varepsilon}{3\alpha\varepsilon}.$ (18)
To proceed, it is interesting to analyze the consistency between our results and the usual form of the AdS/CFT correspondence. Moreover, in the present analysis, we shall consider that there exists an upper bound to the exponent $\alpha$ in Eq (9) to obtain a saturation limit for $C_{h}$, according to its second law.
In this sense, we note from Eqs. (10) and (15), that
$L\sim N^{\frac{2\alpha\varepsilon-3}{6\varepsilon}}.$ (19)
To have consistency with the AdS/CFT correspondence, $L$ must depend on a positive power of $N$ in the sense of obtaining a low curvature regime when $N$ is large [71]. In this case, we must have
$02\alpha\varepsilon-3>$ 0 . (20)
The only possible solution to $\gamma$ consistent with the existence of an upper bound to the exponent $\alpha$ is
$\gamma=\frac{3+3\varepsilon-4\alpha\varepsilon}{3\alpha\varepsilon}\geq 0.$ (21)
In fact, such a result together with Eq. (20) provides:
$\varepsilon$ $>$ $1\;;$
$\frac{3}{2\varepsilon}<$ $\alpha$ $\leq\frac{3(\varepsilon+1)}{4\varepsilon},$ (22)
where the upperbound to $\alpha$ is given by
$\alpha=\frac{3(\varepsilon+1)}{4\varepsilon}.$ (23)
In this way, according to Eqs. (17) and (21), $H_{0}$ must depend on a non-negative power of the holographic complexity. It corresponds to the central result of our work, which can introduce a new understanding of the very nature of $H_{0}$ based on quantum gravity: that $H_{0}$ must be related to the amount of quantum complexity required for cosmological spacetime comes to the existence.
Considering the second law of holographic complexity, $H_{0}$ must increase during the evolution of the cosmos till it saturates at a maximal value. It must shed some light on the issue of Hubble tension by tracing a way $H_{0}$ can have different values throughout the universe’s evolution.
6 Conclusions and discussions.
In the present paper, we have used the quantum gravity scenario introduced in [20, 21] to approach the important issue of Hubble tension. In this case, it was possible to trace an interesting relationship between $H_{0}$ and the holographic complexity associated with the emergence of cosmological spacetime.
As an important consequence, we have obtained a link between $H_{0}$ and the second law of quantum complexity. According to such a law, quantum complexity must increase with time till it saturates at a maximal value [22, 68, 69].
In this case, from the results found out in the present paper, $H_{0}$ must increase with the universe’s holographic complexity during the evolution of the cosmos till it saturates. At this point, such a parameter must reach a constant behavior, recovering the description given by the standard cosmological model. This is in line with several results that have pointed out the possibility that $H_{0}$ must vary throughout the cosmological evolution [17, 18, 19, 78, 79].
In front of such results, the difference between the values obtained to $H_{0}$ when one considers early and late-time measurements must be only a testimony that our universe has evolved from a relatively simple state at its early times to a very complex place containing structures like galaxies, stars, quasars, and black holes. In this sense, it is interesting that black holes have been pointed as the structures that carry the maximal amount of quantum complexity in nature [80].
Such a quantum gravitational reasoning about the Hubble tension can bring us a change of perspective on cosmological observations by considering a quantum gravitational standpoint. In this case, it is interesting to note that, in the context of the theory introduced in [20, 21], there is no role for observers (quantum reference frames) in a quantum gravitational context. This is because quantum reference frames are diluted among the fundamental degrees of freedom of the theory since there is no fundamental distinction between them and quantum correlations in this scenario.
Thus an important change of perspective on the description of the macrocosmos becomes necessary in front of the results found in the present work. One can catch sight of this by noting that $C_{h}^{-\gamma}\sim H_{0}^{-1}$ gives a distance scale in cosmology, corresponding to the observable universe’s radius. Consequently, it is possible to understand the different $H_{0}$ values as a manifestation of the universe’s complexity dependence on distance. In this case, the universe’s complexity must decrease when we consider large distance scales, which can be understood on the basis of an important detail of the theory introduced in the present paper: the spin network description of the microscopic structure of spacetime.
In this way, according to Eqs. (9) and (11), the amount of quantum complexity necessary to build spacetime is given in terms of the number of area quanta patched on the universe’s causal diamond horizon by the spin network legs. In this scenario, the number of quanta in a quantized spacetime must depend on what we observe. Particularly, on the scale of distance we observe, where the number of area quanta must decrease as we increase the distance scale due to coarse-graning effects [81, 82].
Therefore, the results found in the present paper can offer not only an understanding of the important issue of the Hubble tension based on quantum gravity but also a new perspective on cosmological observations through a possible link between the micro and the macrostructure of the cosmos: the less holographic complexity we observe, the further into the universe we can see.
In this way, different observers in the universe must correspond only to different measures of the holographic complexity associated with the emergence of spacetime by considering different regions of the cosmos. It can open up the range of investigation in quantum cosmology by revealing a new perspective of the world where quantum complexity assumes a central role.
Future explorations in Astrophysics can be done in this direction by considering, e.g., gravitational waves [83], as future detectors become able to measure the quantum complexity amount carried by this kind of signal.